# Language-Integrated Recursive Queries

### v1

@@ Abstract.
Performance-critical industrial applications, including large-scale program, network, and distributed system analyses, rely on fixed-point computations. The introduction of recursive common table expressions (CTEs) using the WITH RECURSIVE keyword in SQL:1999 extended the ability of relational database systems to handle fixed-point computations, unlocking significant performance advantages by allowing computation to move closer to the data. Yet with recursion, SQL becomes a Turing-complete programming language and, with that, unrecoverable safety and correctness risks.
SQL itself lacks a fixed semantics, as the SQL specification is written in natural language, full of ambiguities that database vendors resolve in divergent ways. As a result, reasoning about the correctness of recursive SQL programs must rely on isolated mathematical properties of queries rather than wrestling a unified formal model out of a language with notoriously inconsistent semantics. To address these challenges, we propose a calculus, , that automatically derives mathematical properties from embedded recursive queries and, depending on the database backend, rejects queries that may lead to the three classes of recursive query errors — database errors, incorrect results, and non-termination. We introduce TyQL, a practical implementation in Scala for safe, recursive language-integrated query. Using Named-Tuples and type-level pattern matching, TyQL ensures query portability and safety, showing no performance penalty compared to raw SQL strings while unlocking a three-orders-of-magnitude speedup over non-recursive SQL queries.
@@ 1. Introduction
Relational data management systems (RDBMS) are the most popular option for efficient data storage and analytics. In SQL:1999, support was added for recursive common table expressions (CTEs) using the WITH RECURSIVE keyword, enabling fixed-point workloads to execute within RDBMS. Recursive SQL unlocks significant performance gain by pushing computation closer to the data: without WITH RECURSIVE, users must either iterate within the application, use procedural extensions to SQL, or rely on domain-specific backends such as Datalog engines. Repeated round-trips between the application and database or context switches between procedural and plain SQL have significant overhead compared to a single recursive query (Hirn and Grust, 2021). Ideally, applications with fixed-point computations would be able to take advantage of the billions of dollars invested in optimizing relational systems, accessing cutting-edge research and engineering advances without needing to be retrofitted into specialized, domain-specific systems. However, the use of recursion within SQL is widely regarded as “powerful and versatile but notoriously hard to grasp and master” (Hirn and Grust, 2023).
@@
SQL does not have well-established formal semantics as the SQL specification is written in natural language, and is inconsistently followed across RDBMS implementations. Each database vendor adopts its own interpretation, leading to differences in evaluation order, type coercion, null handling, and even fundamental query behavior. As a result, there is no single type-safe or semantically well-defined compilation target for SQL—only a patchwork of best-effort approximations, none widely applied industrially. Yet SQL remains the most popular declarative language due to its powerful optimizability.
Language-integrated query addresses this problem from within the programming language by embedding query constructs using comprehension syntax that supports operations over collections analogous to SQL operations over tables.
Language-integrated query provides safety and portability by allowing users to write their query once and have the language compiler specialize it for multiple backends. The most successful language-integrated query library is the LINQ framework for .NET (Meijer et al., 2006), but the concepts have been applied to many popular languages. Figure 1 shows a classic database query in SQL; canonically in Scala, one of the go-to languages for big data processing and analytics; and our language-integrated query library for Scala, TyQL.
| ⬇ CREATE TABLE Product ( id PRIMARY KEY, name TEXT, price NUMERIC); CREATE TABLE ShipInfo ( id PRIMARY KEY, date DATE); CREATE TABLE Purchase ( id PRIMARY KEY, sid INT, pid INT, cnt INT, total NUMERIC); SELECT pr.name AS name, pu.id AS id, si.date AS sd FROM Purchase pu, Product pr, ShipInfo si WHERE pu.pid = pr.id AND pu.sid = si.id | ⬇ case class Product( id: Int, name: String, price: Double) case class ShipInfo( id: Int, date: LocalDate) case class Purchase( id: Int, sid: Int, pid: Int, cnt: Int, total: Double) def tables = ( prods = Seq[Product](), sInfos = Seq[ShipInfo](), purchs = Seq[Purchase]() ) for purch <- tables.purchs prod <- tables.prods si <- tables.sInfos if purch.pid == prod.id && purch.sid == si.id yield (name = prod.name, id = purch.id, sd = si.date) | ⬇ case class Product( id: Int, name: String, price: Double) case class ShipInfo( id: Int, date: LocalDate) case class Purchase( id: Int, sid: Int, pid: Int, cnt: Int, total: Double) def tables = ( prods = Table[Product](), sInfos = Table[ShipInfo](), purchs = Table[Purchase]() ) for purch <- tables.purchs prod <- tables.prods si <- tables.sInfos if purch.pid == prod.id && purch.sid == si.id yield (name = prod.name, id = purch.id, sd = si.date) |
Classic query comparison.
@@
However, with the addition of WITH RECURSIVE, SQL becomes a Turing-complete programming language and with that comes a host of new correctness and safety concerns orthogonal to classic language-integrated query problems like nested queries or data-type safety. Programming languages have distinct recursion semantics that may only partially align with the SQL specification — and even less with the practical implementations found across modern database systems. The support and implementation of recursive queries vary more widely compared to more well-established SQL features (Winand, 2024). As shown in Figure 2, recursive SQL queries can return incomplete results, throw runtime errors, or never terminate as databases do not effectively identify and prevent the semantic properties of queries that cause problematic behavior. The most reliable way to reason about SQL does not come from a unified semantics imposed across implementations, but rather from the independently applicable and composable mathematical properties of fixed-point computations that underlie relational query evaluation. We propose a solution based on language-integrated query that constrains recursive queries at compile-time using classic mathematical properties, generating only the safe subset of SQL that will not trigger problematic behavior.
To the best of our knowledge, this is the first work that considers the problem of safe language-integrated recursive SQL. We propose holistic compilation of embedded recursive queries to SQL’s WITH RECURSIVE to push computation to the database and enable significant performance benefits. We design the type system to automatically derive relevant mathematical properties of recursive language-integrated queries and reject unsafe queries that risk causing unwanted database behavior. This paper makes the following contributions:
Previous work in language-integrated query did not support recursive SQL. With the addition of a fixed-point operator, SQL becomes a Turing-complete language, leading to new and complex challenges to query correctness and safety. In Section 2.2 we identify the 3 classes of ways that recursive queries “go wrong”: runtime exception, incorrect results, and nontermination, and show what mathematical properties of queries are responsible for each behavior.
We present a calculus, , that automatically checks the mathematical properties responsible for these classes of errors at the host-language compile-time. Sections 4 and 5 show how uses type classes, linear types, and union types to independently encode constraints for each targeted mathematical property so that the generated query is specialized to the semantics of the database backend.
@@
In Section 7 we propose a benchmark of recursive queries adapted from industrial benchmarks and academic works in the domains of recursive SQL, Datalog, and Graph Database systems. We conduct a survey of modern relational database systems with respect to query behavior.
@@
We then evaluate TyQL with regard to query coverage and the performance of TyQL with alternative approaches, showing no performance penalty compared to raw SQL strings and a three-orders-of-magnitude speedup over state-of-the-art language-integrated query libraries using non-recursive SQL queries with an in-memory embedded database.
2. Background on Recursive Querying
@@
The algorithms driving modern applications demand increasingly sophisticated query capabilities beyond the limits of traditional select-project-join-aggregate queries. Performance-critical industrial applications, including large-scale program, network, and distributed system analyses, rely on fixed-point computations for data analysis.
2.1. Recursive Query Languages
Datalog is a rule-based declarative query language based in logic programming that has been widely used for recursive query processing, particularly program analysis (Scholz et al., 2016). Datalog programs take the form of a finite set of facts and rules. Datalog rules take the following form:
@@ 2.1. Recursive Query Languages
are predicate symbols, also called relations, and and are terms that can be either variables or constants. The bottom-up evaluation of a Datalog program repeatedly executes unions of conjunctive queries for each derived predicate symbol until an iteration passes where no new facts are discovered (Gottlob et al., 1989). The Semi-Naive evaluation technique only uses facts from the previous iteration to discover new facts (Ullman, 1989).
WITH RECURSIVE was officially added to the SQL standard in 1999 (International Organization for Standardization (1999), ISO). Prior, hierarchical or recursive relationships (for example organizational charts, family trees, graph data, etc.) required use of Datalog, application-side iteration on data extracted from the database, or specialized, non-standard SQL extensions, for example Oracle’s CONNECT BY or PL/SQL. Even after WITH RECURSIVE became standardized, databases have added more powerful or expressive support for recursive queries. For example, DuckDB (duckdb, 2019) allows mutual recursion and MariaDB supports mutual and non-linear recursion (Shalygina and Novikov, 2017). RaSQL (Wang et al., 2020) extends recursive queries to allow certain non-monotone aggregations. Datalog engines, for example Soufflé (Scholz et al., 2016), support expressive forms of recursion like mutual or non-linear recursion but do not implement the SQL standard.
| Recursion | Nonlinear | Mutual | CYCLE | UNION [distinct] | |
| MySQL | ✓(2017) | • (2017) | ✗ | ✗ | ✓ (2017) |
| OracleDB | ✓ (2009) | ✗ | ✗ | ✓ (2009) | ✗ |
| PostgreSQL | ✓ (2009) | ✗ | ✗ | ✓ (2021) | ✓ (2009) |
| SQL Server | ✓ (2005) | • (2008) | ✗ | ✗ | ✗ |
| SQLite | ✓ (2014) | • (2020) | ✗ | ✗ | ✓ (2014) |
| MariaDB | ✓ (2017) | ✓ (2017) | ✓ (2017) | ✓(2020) | ✓ (2017) |
| DuckDB | ✓ (2020) | • (2020) | • (2020) | ✗ (2020) | ✓ (2020) |
@@
Table 1 shows documented recursion support across several modern relational database systems. ✗ indicates that the database does not support the feature and will throw a relevant error message. • indicates syntactic support for a feature, namely, the query will not throw an error but the system may not necessarily have the requisite internal implementation to execute the query successfully. It is possible to run non-linear recursive queries on some databases using syntactic tricks to obfuscate that the query is not linear, e.g. PostgreSQL, but it is not officially supported. ✓ indicates documented support. As illustrated in Table 1, the technical landscape is wide and ever-growing, and no two modern databases support the same set of features. As a result, constraints on recursive queries must be applied independently and specialized to the database.
2.2. How do recursive queries “go wrong”?
@@
In this section, we classify the problems that arise with recursive queries into three areas and identify the mathematical properties responsible for each class of error. The three emergent behaviors targeted are (1) runtime exceptions thrown by the database, (2) incomplete results returned from a query, and (3) nontermination. We show how each targeted behavior is a result of a particular mathematical property, then discuss the cases in which a user may want to relax these restrictions.
2.2.1. Recursive Query Runtime Exception
It is well-known in the Datalog and recursive query literature that to guarantee that the unique and minimal fixed point will be found in a finite number of steps, operations within recursive queries must be monotonic under the ordering of set inclusion: for a monotonic query and databases and , if then (Bancilhon and Ramakrishnan, 1986). Therefore, all non-monotonic operators, i.e. aggregation or negation, must happen between strata of recursive queries to guarantee termination (that is, aggregations can only be applied to the results of recursive queries after they have completed, not to results within the bodies of recursive queries).
@@ 2.2.1. Recursive Query Runtime Exception
For example, Figure 3 shows the widely-used Bill-of-Materials problem (Wang et al., 2020) that models items sold by a business that are made out of sub-parts. We use the relation SubParts(part, sub) for each item a business sells and its sub-parts (and sub-part sub-parts, etc.); and the relation BasicParts(part, days) for base parts and how many days it takes for that part to arrive from a supplier. The query waitFor determines when a part will be ready, given that a part will be ready on the day the last subpart arrives. Naively, this query can be represented as Figure 3. However, the presence of the MAX aggregation within the body of the recursive query definition means that the query is not monotonic under the ordering of set inclusion. The SQL standard specifies that only monotonic queries are supported. This constraint is usually checked by the query engine, so the emergent behavior is a run-time error for the application as it must wait for the round-trip time for the query to be sent to the database and the error returned. All widely-used commercial databases and Datalog engines officially support only monotonic operations within recursion, so for the query in Figure 3 to pass a database query compiler check, the MAX aggregation must be moved out of the recursive query. The set of database query errors that are unrelated to recursion, for example datatype errors, are well covered by previous work in language-integrated query that is complementary to our approach (Cheney et al., 2013). Therefore, the mathematical property that leads to recursive query runtime exceptions is non-monotonicity.
@@
However, recent advances in recursive query engines have shown that some forms of aggregation (Wang et al., 2020; Abo Khamis et al., 2022) are permissible within recursion without losing termination guarantees. While this has not yet been implemented widely in commercial systems, it shows when a user may still want to “turn off” the monotonicity constraint for certain backends.
2.2.2. Incomplete Results

### v2

@@ Abstract.
Performance-critical applications, including large-scale program analyses, graph analyses, and distributed system analyses, rely on fixed-point computations. The introduction of recursive queries using the WITH RECURSIVE keyword in SQL:1999 extended the ability of relational database systems to handle fixed-point computations, unlocking significant performance advantages by allowing computation to move closer to the data. Yet, with the addition of recursion, SQL becomes a Turing-complete programming language with new correctness and safety risks.
SQL lacks a fixed semantics, as the SQL specification is written in natural language, full of ambiguities that database vendors resolve in divergent ways. As a result, reasoning about the correctness of recursive SQL programs must rely on isolated properties of queries rather than wrestling a unified formal model out of a language with notoriously inconsistent semantics. To address these challenges, we propose a calculus, , that derives properties from embedded recursive queries using the host-language type system and, depending on the database backend, rejects queries that may lead to the three classes of recursive query errors - runtime database exceptions, incorrect results, and non-termination. We introduce TyQL, a practical implementation in Scala for safe, recursive language-integrated query. TyQL uses modern type system features of Scala 3, namely Named-Tuples and type-level pattern matching, to ensure query portability and safety. Furthermore, TyQL shows no performance penalty compared to raw SQL strings while unlocking a three-orders-of-magnitude speedup over non-recursive SQL queries.
@@ 1. Introduction
Fixed-point workloads contain computations that apply a query repeatedly, over the result of the previous iteration, until the result no longer changes. SQL:1999 (International Organization for Standardization, 1999) introduced the WITH RECURSIVE keyword, enabling fixed-point workloads to be executed within relational database management systems (RDBMS). Recursive SQL unlocks significant performance gain, in part by pushing computation closer to the data (Hirn and Grust, 2021). Ideally, applications with fixed-point computations would be able to leverage the billions of dollars invested in optimizing RDBMS, accessing cutting-edge research and engineering advances without needing to be retrofitted into specialized, domain-specific systems, such as Datalog engines.
@@
Despite excellent performance, the use of recursion within SQL is widely regarded as “powerful and versatile but notoriously hard to grasp and master” (Hirn and Grust, 2023). The difficulty is exacerbated by the lack of a well-established formal semantics, as the SQL specification is written in natural language and is inconsistently followed across RDBMS implementations. Each database vendor adopts its own interpretation, leading to differences in evaluation order, type coercion, null handling, and even fundamental query behavior (Neumann and Leis, 2024). As a result, there is no single type-safe or semantically well-defined compilation target for SQL, only a patchwork of best-effort approximations that either make simplifying assumptions or target a limited subset of SQL (Negri et al., 1991; Guagliardo and Libkin, 2017; Veanes et al., 2010; Chu et al., 2017; Malecha et al., 2010; Ricciotti and Cheney, 2022), none widely applied industrially.
Language-integrated query addresses inconsistencies across RDBMS from within the programming language, providing safety and portability by allowing users to write their query once and have the language compiler type-check and specialize it for multiple backends. The most successful language-integrated query library is the LINQ framework for .NET (Meijer et al., 2006), and the concepts have been widely adopted across modern programming languages and databases (slick, 2011; quill, 2016; jslinq, 2025; diesel, 2025). Yet neither LINQ nor other language-integrated libraries support recursion. With the addition of WITH RECURSIVE, SQL becomes a Turing-complete programming language, and with that comes a host of new concerns orthogonal to classic language-integrated query problems like nested queries or data-type safety.
@@
Figure 1 shows how recursive SQL queries can throw runtime errors, return incorrect results, or never terminate. Furthermore, the support and implementation of recursive queries vary more widely across commercial databases compared to other SQL features (Winand, 2024). As a result, the most effective and practical way to reason about recursive SQL does not come from a unified semantics imposed across database implementations.
The key insight behind our paper is to provide a DBMS-agnostic reasoning framework over recursive queries based on independently applicable and composable mathematical properties of fixed-point computations that determine how queries will behave: range-restriction, monotonicity, mutual-recursion, linearity, set-semantics, and constructor-freedom. We propose a solution based on language-integrated query that constrains queries at application compile-time, enabling significant performance benefits while generating only the safe subset of SQL that will not trigger runtime exceptions, nontermination, or incorrect results.
Figure 2 shows a classic recursive database query for the transitive closure of a directed graph in our language-integrated query library for Scala, TyQL, and in SQL. We design the type system to automatically derive the six relevant mathematical properties and, depending on the database backend, reject unsafe queries. This paper makes the following contributions:
Previous work in language-integrated query did not support recursive SQL. With the addition of a fixed-point operator, SQL becomes a Turing-complete programming language, leading to new and complex challenges to query correctness and safety. In Section 2 we provide background on recursive database queries and in Section 3 we identify the three classes of ways that recursive queries “go wrong”: runtime exception, incorrect results, and nontermination, and show the mathematical properties of queries that are responsible for each behavior.
We present a calculus, , that automatically checks the six properties responsible for these classes of errors at the host-language compile-time. Sections 5 and 6 show how uses type classes, linear types, and union types to independently encode constraints for each targeted query property so that the generated query is specialized to the semantics of the database backend.
@@
In Section 8 we propose a benchmark of recursive queries adapted from industrial benchmarks and academic works in the domains of recursive SQL, Datalog, and Graph Database systems. We conduct a survey of modern relational database systems with respect to query behavior. We then evaluate TyQL with regard to query coverage and the performance of TyQL with alternative approaches, showing no performance penalty compared to raw SQL strings and a three-orders-of-magnitude speedup over state-of-the-art language-integrated query libraries using non-recursive SQL queries with an in-memory embedded database.
Examples of Transitive Closure Query in Recursive SQL, and TyQL.
2. Background on Recursion in Databases
@@
Modern applications demand increasingly sophisticated query capabilities beyond the limits of traditional select-project-join-aggregate queries. Performance-critical industrial applications, including large-scale program analyses, network analyses, artificial intelligence, and distributed system analyses, rely on fixed-point computations (Sahebolamri et al., 2022; Backes et al., 2019; Das et al., 2019; Darwiche, 2020; Bellomarini et al., 2018; Vistrup, 2025).
@@
Datalog (Gottlob et al., 1989) is a recursive query language based on logic programming that has been particularly useful for program analysis (Scholz et al., 2016). A Datalog program is a finite set of facts and rules where rules take the form: . are predicate symbols, also called relations, and and are terms that can be either variables or constants. The bottom-up evaluation of a Datalog program repeatedly executes unions of conjunctive queries for each derived predicate symbol until an iteration passes where no new facts are discovered (Gottlob et al., 1989). The Semi-Naive evaluation technique only uses facts from the previous iteration to discover new facts (Ullman, 1989).
WITH RECURSIVE was added to the SQL standard in 1999 (International Organization for Standardization, 1999), based on the same theoretical foundation and Semi-Naive evaluation technique developed for Datalog. Prior, hierarchical or recursive relationships (e.g., organizational charts, family trees, graph data, etc.) required use of Datalog, application-side iteration on data extracted from the database, or specialized, non-standard SQL extensions, for example Oracle’s CONNECT BY keyword or procedural extensions like PL/SQL. It has been shown that repeated round-trips between the application and database or context switches between procedural and plain SQL have significant overhead compared to a single recursive query (Hirn and Grust, 2021). Recursive queries define recursive relations, and are composed of a base case query and a recursive case query that references the recursive relations. For example, the query in Figure 2 defines a recursive relation named ”path”, the base case is the ”edges” relation, and the recursive case joins the ”edges” and ”path” relations, producing the transitive closure of all nodes.
@@
Since 1999, databases have added even more powerful or expressive support for recursion. Table 1 shows documented recursion support across several modern relational database systems. ✗ indicates that the database does not support the feature and will throw a relevant error message. • indicates syntactic support for a feature, namely, the query will not throw an error but the system may not necessarily have the requisite internal implementation to execute the query successfully.
| WITH | Range-Rest | Agg | Mutual | Non-linear | Set | Const-Free | |
| RECURSIVE | (P1) | (P2) | (P3) | (P4) | (P5) | (P6) | |
| MySQL | ✓ (2017) | ✓ (2017) | ✗ | ✗ | • (2017) | ✓ (2017) | ✓ (2017) |
| OracleDB | ✓ (2009) | ✓ (2009) | ✗ | ✗ | ✗ | ✗ | ✓ (2009) |
| PostgreSQL | ✓ (2009) | ✓ (2009) | ✗ | ✗ | ✗ | ✓ (2009) | ✓ (2009) |
| SQL Server | ✓ (2005) | ✓ (2005) | ✗ | ✗ | • (2008) | ✗ | ✓ (2005) |
| SQLite | ✓ (2014) | ✓ (2014) | ✗ | ✗ | • (2020) | ✓ (2014) | ✓ (2014) |
| MariaDB | ✓ (2017) | ✓ (2017) | ✗ | ✓ (2017) | ✓ (2017) | ✓ (2017) | ✓ (2017) |
| DuckDB | ✓ (2020) | ✓ (2020) | ✓ (2020) | • (2020) | • (2020) | ✓ (2020) | ✓ (2020) |
Range-restriction. The “Range-Rest” column in Table 1 indicates queries must be range-restricted. The Datalog literature defines the property of range-restriction (Gottlob et al., 1989) as the constraint that variables on the left-hand-side of a Datalog rule must appear in the right-hand-side. Applied to recursive SQL queries, this constraint requires the project clause to contain only constants or references to columns present in relations in the FROM clause. All RDBMS require range-restriction.
Monotonicity. The “Agg” column indicates support for aggregation operations within the bodies of recursive queries. So far, the only commercial database to allow for aggregations within recursive queries is DuckDB, but academic systems like RaSQL (Wang et al., 2020) extend recursive queries to allow certain restricted aggregations. Aggregations are restricted because it is well-known in the Datalog and recursive query literature that to guarantee that the unique and minimal fixed point will be found in a finite number of steps, operations within recursive queries must be monotonic under the ordering of set inclusion. A monotonic query is defined as a query and databases and , such that if then (Bancilhon and Ramakrishnan, 1986). A query is monotonic if adding more data to its input does not remove data from its output, and aggregations like MAX or NOT EXISTS can violate this property. Datalog variants with aggregation generally require stratification, where recursive relations are grouped into layers, or strata, where each layer can only depend on the results of earlier layers. To ensure termination, aggregation is allowed across strata, but not within a single stratum. Recursive SQL enforces stratification by construction, as each WITH RECURSIVE query is defined as one stratum. Because aggregation is allowed between recursive queries, all the commercial RDBMS presented in Table 1 except DuckDB support stratified aggregation, while DuckDB is more expressive than stratified Datalog.
Mutual-recursion. The “Mutual” column indicates support for mutually recursive queries, where two or more recursive relations refer to each other in a cyclic dependency. Support for mutual recursion is essential for expressing many static analyses or bidirectional graph traversals, and is built into Datalog semantics but not SQL. Recently, MariaDB added support for mutual recursion, and DuckDB inlines relations such that it is possible to express some mutually recursive queries.
Linearity. The “Non-linear” column indicates support for non-linear recursive queries, in which a recursive relation is referenced more than once within a recursive subquery. Non-linear queries are particularly useful for program analysis queries (Wang et al., 2017; Madsen et al., 2016; Abeysinghe et al., 2024). DuckDB does not check queries for linearity, PostgreSQL applies a simple and easily avoided syntactic check, and MariaDB supports non-linear queries.
Set-semantics. The “Set” column indicates support for the UNION operator to combine the base and recursive cases, which uses set semantics. Without UNION, the default is UNION ALL, which uses bag (i.e., multiset) semantics. Some RDBMS, e.g., OracleDB, require the use of bag semantics.
Constructor-freedom. The “Const-Free” column indicates support for queries that violate the constructor-freedom property. The Datalog literature refers to queries that do not contain interpreted functions over infinite domains (e.g., integer arithmetic) as having the constructor-freedom property (Arntzenius and Krishnaswami, 2016). All RDBMS support constructors.
Pure Datalog does not support constructors or aggregations. Modern Datalog engines, for example Soufflé (Scholz et al., 2016), support expressive forms of recursion like mutual or non-linear recursion and use set semantics but miss out on the wide adoption and significant engineering resources that go into making commercial RDBMS efficient. As illustrated in Table 1, the technical landscape is wide and ever-growing, and no two modern databases support the same set of features. The lack of alignment among commercial RDBMS renders a single, unified formal semantics untenable as a foundation for correctness checking of recursive queries; practical systems must instead support a flexible and composable framework that adapts to the specific capabilities of each backend.
3. How do recursive queries “go wrong”?
@@
In this section, we classify the problems that arise with recursive queries into three areas based on the emergent database behavior and define the mathematical query properties responsible for each class of error. As illustrated in Figure 1, the three behaviors targeted are runtime exception (B1), incorrect results (B2), and nontermination (B3). We define six mathematical query properties: range-restriction (P1), monotonicity (P2), mutual-recursion (P3), linearity (P4), set-semantics (P5), and constructor-freedom (P6), and show how each behavior B1-B3 may result from violating one or more of P1-P6, providing an example query (shown in Figure 3) for each query Q1-Q3 in Figure 1. Lastly, we discuss the cases where a user may want to selectively apply or relax the restriction of each property, even if the RDBMS supports it.
3.1. Recursive Query Runtime Exception (B1)
@@
Language-integrated query targets many problems associated with query writing: datatype or schema errors, for example, falsely assuming a table to have a particular column name or data type; syntactic errors like simple typos; security vulnerabilities such as SQL injection; and structural mistakes like HAVING without GROUP BY. Without language integration, the RDBMS query compiler will identify these errors and throw an exception, which will be a runtime error for the application. These types of errors are already well-covered by existing language-integrated query techniques that are complementary to our approach (Cheney et al., 2013). However RDBMS query compilers throw recursion-specific exceptions that are not caught by existing techniques.
@@
Example. The SQL standard specifies that only monotonic queries are supported. Practically, this means that all non-monotonic operators, i.e., aggregation or negation, must happen between recursive SQL queries to guarantee termination. For example, Figure 3 shows the widely-used Bill-of-Materials problem (Wang et al., 2020) that models items sold by a business that are made out of sub-parts. We use the relation SubParts(part, sub) for each item a business sells and its sub-parts (and sub-sub-parts, etc.); and the relation BasicParts(part, days) for base parts and how many days it takes for that part to arrive from a supplier. The query waitFor determines when a part will be ready, given that a part will be ready on the day the last subpart arrives. However, the presence of the MAX aggregation within the body of the recursive query definition means that the query is not monotonic under the ordering of set inclusion.
This constraint is usually checked by the query compiler, so the emergent behavior is a runtime error for the application as it must wait for the round-trip time for the query to be sent to the database and the error returned. Most widely-used commercial RDBMS and Datalog engines officially support only monotonic operations within recursion, so for the query in Figure 3 to pass a database query compiler check, the MAX aggregation must be moved out of the recursive query. Therefore, it is the violation of property P2: monotonicity that leads to behavior B1: recursive query runtime exception on the systems that do not support non-monotonic recursion. Queries that violate P1: range-restriction will also always cause runtime exceptions. Additionally, many RDBMS do not support mutual or non-linear recursion, so queries that violate property P3: mutual-recursion or P4: linearity will also result in a runtime error for those systems. In Figure 1, queries that violate P1, P2, P3, or P4 and are checked by the query compiler are represented by query Q3.
@@
However, recent advances in recursive query engines have shown that some forms of aggregation (Wang et al., 2020; Abo Khamis et al., 2022) are permissible within recursion without losing termination guarantees. While this has not yet been implemented widely in commercial systems, it shows when a user may still want to “turn off” the monotonicity constraint for certain backends without risking nontermination.
3.2. Incorrect Results (B2)
A linearly recursive function is a recursive function that contains exactly one recursive call. Similarly, the SQL standard defines a linear query to be a query that references the recursive relation once. Around the time that SQL’99 was preparing to add recursive queries to SQL, there was a belief that “most ‘real life’ recursive queries are indeed linear” (Bancilhon and Ramakrishnan, 1986). While this belief is no longer widely held, this assumption is built into the implementation of many RDBMS.